OpenStudy (darkigloo):

Determine whether the sequence a(n)=arctan(20n^18) converges or diverges.

OpenStudy (darkigloo):

if it converges find the limit

OpenStudy (anonymous):

i may be wrong. \[\tan \frac{ \pi }{ 2 }\rightarrow \infty\] \[\tan^{-1} \infty =\frac{ \pi }{ 2 }\] \[a _{n}=\frac{ \pi }{ 2 },as~n \rightarrow \infty \] or an not approaches 0 as n approaches infinity so an diverges.

OpenStudy (darkigloo):

ok thank you :)

OpenStudy (anonymous):

yw